# Dark matter admixed neutron stars with a realistic nuclear equation of state from chiral nuclear interactions

### v1

@@ Abstract
We study the effects of dark matter on the structural properties of neutron stars. In particular we investigate how the presence of a dark matter component influences the mass-radius relation, the value of the maximum mass of a neutron star and others stellar properties. To model ordinary matter we use a state-of-the-art equation of state of $\beta$-stable nuclear matter obtained using the Brueckner-Hartree-Fock quantum many-body approach starting from two-body and three-body nuclear interactions derived from chiral effective field theory. The dark matter component of the star is modeled as a non-self-annihilating system of spin $1/2$ fermions and its equation of state as an ideal relativistic Fermi gas. The equilibrium configurations of these dark matter admixed neutron stars (DANS) are calculated by solving a generalization of the Tolman-Oppenheimer-Volkoff equations to the case where the system consists of two perfect fluids interacting solely through gravity. We find that, depending on the dark matter particle mass $m_{\chi}$, one can have somehow opposite effects on the stellar properties. In the case $m_{\chi}=1\,\mathrm{GeV}$, the stellar gravitational maximum mass $M_{max}$ decreases, whereas in the case $m_{\chi}=0.1\,\mathrm{GeV}$, $M_{max}$ increases with respect to the maximum mass of ordinary neutron stars. We also show that the presence of dark matter has indirect sizeable effect on the proton fraction in the ordinary matter fluid and, in the case $m_{\chi}=1\,\mathrm{GeV}$, results in a decrease of the threshold gravitational mass $M_{tot}^{durca}$ for having direct URCA processes and fast stellar cooling. Finally we study the stability of dark matter admixed neutron stars with respect to radial perturbations.
I Introduction
@@
In spite of its many successes, Einstein’s theory of general relativity provides predictions that are at variance with astrophysical observational data on the kinematics of self-gravitating systems as galaxies and clusters of galaxies Corbelli and Salucci (2000); Nesti et al. (2023). In order to reproduce these data using general relativity requires the introduction of a new type of matter, known as dark matter (DM), that is not directly observable except through its gravitational effects on ordinary matter. Cosmological data, currently explained by the $\Lambda$-cold dark matter ($\Lambda$CDM) model Peebles (1982), indicate that DM should be responsible for about 23% percent of the total energy density of the Universe while ordinary matter (OM) contributes only 4%. The remaining portion is due to dark energy responsible for the accelerated expansion of the Universe Goldhaber and Perlmutter (1998); Garnavich et al. (1998).
@@
The main feature of DM is that its cross section for electromagnetic processes must be extremely small or zero. As a result, DM cannot lose energy by emitting electromagnetic radiation, as in the case of ordinary matter, and thus DM does not collapse at the center of galaxies. In fact, the rotation curve profiles of disc galaxies (i.e. the plot of the orbital speeds of visible stars or gas in the galaxy versus their radial distance from the galaxy’s centre) are in agreement with the presence of an extended and diffuse halo of DM (see Sofue (2013); Pato et al. (2015); Lin and Li (2019)).
@@
None of the particles of the standard model meet the requirements for explaining the numerous astrophysical observations and cosmological data that require the existence of dark matter. Thus candidate particles to constitute dark matter most often emerge from extensions of the standard model. A much-studied class of such particles are WIMPs (weakly interacting massive particles), and many active experiments are attempting to reveal their existence. No such experiments, however, have so far obtained convincing results on the existence of DM particles even though very stringent constraints on the interaction cross section of dark matter with ordinary matter have been produced (see e.g. Aprile et al. (2017)).
@@
Due to the very high densities ($10^{14}$$10^{15}\,\mathrm{g/cm^{3}}$) in their cores, the extreme surface gravity ($g\sim 10^{12}\,\mathrm{m/s^{2}}$) and their ages (up to $10^{10}\,\mathrm{yr}$), neutron stars are very good candidates where dark matter can be captured and can accumulate in sufficient amounts to influence their structural and evolutionary properties. Neutron stars can thus be used as natural detectors to identify DM and possibly unveil its nature Bertone and Fairbairn (2008); de Lavallaz and Fairbairn (2010); Kouvaris and Tinyakov (2010, 2011); Leung et al. (2011); Ciarcelluti and Sandin (2011); Kouvaris et al. (2013); Kouvaris and Nielsen (2015); Ellis et al. (2018); Ivanytskyi et al. (2020); Del Popolo et al. (2020); Kain (2020, 2021); Shakeri and Karkevandi (2024).
@@
This work is organized as follows. In Section II, we introduce the two-fluids Tolman-Oppenheimer-Volkoff (TOV) stellar structure equations. In Section III, we discuss the equation of state for OM and for DM used in the present work. In Section IV we present results obtained by solving the two-fluids TOV equation and discuss the properties of DANS. In Section V we use observational data for properties of neutron stars to determine an upper limit to the dark matter particle mass. In Section VI, we study the stability of DANS equilibrium configurations with respect to small radial perturbations. Finally, in Section VII we give a brief summary of our work.
II Structure equations for DANS
Since the non-gravitational interaction between dark matter and ordinary matter is extremely small (e.g. Bertone et al. (2005); Aprile et al. (2017)), it is possible to split the total energy-momentum tensor as the sum of the energy-momentum tensor of each of the two fluids (OM and DM) and to have covariant conservation for both of them. Accordingly, the equation of state of OM is independent on the state variables of DM and vice versa. In addition, it is assumed that each of the two fluids be a perfect fluid. Based on these assumptions, and further assuming a spherically symmetric and stationary distribution of OM and DM, the stellar structure equations in general relativity for DANS take the following form (see e.g. Kain (2020)), which generalizes the TOV equations to the case of two fluids interacting exclusively through gravity:
$\frac{dP_{j}}{dr}$ $=$ $-G\,\frac{m_{tot}(r)\,\varepsilon_{j}(r)}{c^{2}r^{2}}\,\Bigg{(}1+\frac{P_{j}(r)}{\varepsilon_{j}(r)}\Bigg{)}$ (1)
[…]
@@ II Structure equations for DANS
II.1 Numerical integration
To solve the stellar structure equations (1) and (2) we need to specify the equation of state for the two fluids (see next section) and the appropriate boundary conditions at the center ($r=0$) and at the surface ($r=R_{j}$, $j=OM,DM$) of the matter distribution for each fluid Kain (2020):
@@ II.1 Numerical integration
III Equation of state
III.1 The equation of state of ordinary matter
@@
In this work we model the OM fluid of a DANS as a uniform electric-charge-neutral fluid of neutrons, protons, electrons, and muons. Recently a new microscopic EOS for this system has been obtained in Ref. Bombaci and Logoteta (2018) (hereafter the BL EOS) for the zero temperature case, using the Brueckner-Hartree-Fock (BHF) quantum many-body approach (see Bombaci and Logoteta (2018) and references therein) starting from modern two-body and three-body nuclear interactions derived within chiral effective field theory (ChEFT) (e.g. Machleidt and Entem (2011); Hammer et al. (2020)). These chiral nuclear interactions reproduce with high accuracy the nucleon-nucleon (NN) scattering data and the experimental binding energies of light (A = 3, 4) atomic nuclei Logoteta et al. (2016)).
@@
The BL EOS reproduces the empirical saturation point (i.e. saturation density $n_{0}=0.16\pm 0.01~{}{\rm fm}^{-3}$, and energy per nucleon $E/A|_{n_{0}}=-16.0\pm 1.0~{}{\rm MeV}$) of symmetric nuclear matter, and other empirical properties (symmetry energy $E_{sym}$ and its slope parameter $L$, incompressibility) of nuclear matter at the saturation density $n_{0}$ (see Tab. 2 in Bombaci and Logoteta (2018)). In particular the calculated $E_{sym}(n_{0})$ and $L$ for this EOS fulfill the bounds imposed by the unitary Fermi gas limit Tews et al. (2017); Kievsky et al. (2018). As has been shown in ref. Tews et al. (2017) several EOS models currently used in neutron star structure calculations, in supernova explosions and binary neutron star mergers simulations violate the unitary Fermi gas bounds.
Further the BL EOS is consistent (see figure 2 in Bombaci and Logoteta (2018)) with the measured elliptic flow of matter in heavy-ion collisions experiments Danielewicz et al. (2002).
@@
At supranuclear densities ($n>n_{0}$) the BL EOS model predicts a symmetry energy $E_{sym}(n)$ which is in very good agreement with the empirical constraints derived in Danielewicz and Lee (2014) using the excitation energies of isobaric analog states in nuclei and with additional constraints from neutron skin thickness of heavy nuclei Roca-Maza et al. (2013). As is well known, the symmetry energy $E_{sym}(n)$, and particularly its density dependence is a crucial ingredient to determine the proton fraction in $\beta$-stable nuclear matter Bombaci and Lombardo (1991); Li et al. (2021); Zhang and Li (2021); Bombaci (2022) and ultimately it has an impact on the value of stellar radius and on the thickness of the crust Lattimer and Prakash (2001); Steiner (2006) and on the possibility of having direct URCA processes Lattimer et al. (1991) (see Section IV.2) and thus rapid cooling of neutron stars Page et al. (2004).
@@
When computing static ordinary neutron star configurations the BL EOS (for the $\beta$-stable case) gives Bombaci and Logoteta (2018) a maximum mass $\rm{M}_{\rm max}=2.08\,\mathrm{M_{\odot}}$, with a corresponding central density $\rho_{c}=2.74\times 10^{15}\,\mathrm{g/cm^{3}}$ and radius $\rm{R}(M_{\rm max})=10.26\,\mathrm{km}$ and a quadrupolar tidal polarizability coefficient $\Lambda_{1.4}=385$ (for the $1.4\,\mathrm{M_{\odot}}$ neutron star Logoteta and Bombaci (2019)) compatible with the constraints derived from GW170817 Abbott et al. (2017).
Recently, the BL EOS has been extended Logoteta et al. (2021) to finite temperature and to arbitrary proton fractions. This finite-temperature EOS model has been applied to numerical simulations of binary neutron star mergers Bernuzzi et al. (2020); Endrizzi et al. (2018); Prakash et al. (2021).
Finally, to model the (ordinary) neutron star crust (i.e. for nucleonic density $\leq 0.08$ fm-3) we have used the Baym–Pethick–Sutherland (Baym et al., 1971) and the Negele–Vautherin Negele and Vautherin (1973) EOS.
III.2 The equation of state of dark matter
@@
In the present work we consider non-self-annihilating fermionic DM. This so called fermionic asymmetric dark matter (ADM) Kaplan et al. (2009); Zurek (2014) carry a conserved charge which is analogous to the baryon number in the case of ordinary matter. We describe the ADM fluid as a non-interacting (ideal) gas of fermions with mass $m_{\chi}$ and spin $1/2$. The corresponding EOS is well known (see e.g. Shapiro and Teukolsky (1983)) and is given by the following expressions
@@
IV DANS properties
For a fixed values of the dark matter particle mass $m_{\chi}$, a solution the two-fluid TOV equations (1) and (2) is uniquely identified once we fix the central mass densities (central energy densities) $\rho_{cOM}$ and $\rho_{cDM}$ for the two fluids.
Figure 1: Mass-density profiles of ordinary matter (red) and dark matter (black) to the two-fluids TOV equation for $m_{\chi}=1\ \mathrm{GeV}$, $\rho_{cOM}=2\times 10^{15}\ \mathrm{g/cm^{3}}$ and different values of the dark matter central densities $\rho_{cDM}$. Each line-style correspond to a different choice of $\rho_{cDM}$: $(\rho_{cOM},0)$ (green continuous line), $(\rho_{cOM},10^{15}\ \mathrm{g/cm^{3}})$ (dotted lines), $(\rho_{cOM},4\times 10^{15}\ \mathrm{g/cm^{3}})$ (dash-dot lines), $(\rho_{cOM},2\times 10^{16}\ \mathrm{g/cm^{3}})$ (dashed lines).

### v2

@@ Abstract
We study the effects of dark matter on the structural properties of neutron stars. In particular we investigate how the presence of a dark matter component influences the mass-radius relation, the value of the maximum mass of a neutron star and other stellar properties. To model ordinary matter we use a state-of-the-art equation of state of $\beta$-stable nuclear matter obtained using the Brueckner-Hartree-Fock quantum many-body approach starting from two-body and three-body nuclear interactions derived from chiral effective field theory. The dark matter component of the star is modeled as a non-self-annihilating system of spin $1/2$ fermions at zero temperature and its equation of state as an ideal relativistic Fermi gas. The equilibrium configurations of these dark matter admixed neutron stars (DANS) are calculated by solving a generalization of the Tolman-Oppenheimer-Volkoff equations to the case where the system consists of two perfect fluids interacting solely through gravity. We find that, depending on the dark matter particle mass $m_{\chi}$, one can have somehow opposite effects on the stellar properties. In the case $m_{\chi}=1\,\mathrm{GeV}$, the stellar gravitational maximum mass $M_{max}$ decreases, whereas in the case $m_{\chi}=0.1\,\mathrm{GeV}$, $M_{max}$ increases with respect to the maximum mass of ordinary neutron stars. We also show that the presence of dark matter has indirect sizable effect on the proton fraction in the ordinary matter fluid and, in the case $m_{\chi}=1\,\mathrm{GeV}$, results in a decrease of the threshold gravitational mass $M_{tot}^{durca}$ for having direct URCA processes and fast stellar cooling. Finally we study the stability of dark matter admixed neutron stars with respect to radial perturbations.
keywords:
Neutron Stars , Dark Matter , Equation of State
[…]
1 Introduction
@@
In spite of its many successes, Einstein’s theory of general relativity provides predictions that are at variance with astrophysical observational data on the kinematics of self-gravitating systems as galaxies and clusters of galaxies (Corbelli and Salucci, 2000; Nesti et al., 2023). In order to reproduce these data using general relativity requires the introduction of a new type of matter, known as dark matter (DM), that is not directly observable except through its gravitational effects on ordinary matter. Cosmological data, currently explained by the $\Lambda$-cold dark matter ($\Lambda$CDM) model (Peebles, 1982), indicate that DM should be responsible for about 23% percent of the total energy density of the Universe while ordinary matter (OM) contributes only 4%. The remaining portion is due to dark energy responsible for the accelerated expansion of the Universe (Goldhaber and Perlmutter, 1998; Garnavich et al., 1998).
@@
The main feature of DM is that its cross section for electromagnetic processes must be extremely small or zero. As a result, DM cannot lose energy by emitting electromagnetic radiation, as in the case of ordinary matter, and thus DM does not collapse at the center of galaxies. In fact, the rotation curve profiles of disc galaxies (i.e. the plot of the orbital speeds of visible stars or gas in the galaxy versus their radial distance from the galaxy’s center) are in agreement with the presence of an extended and diffuse halo of DM (see (Sofue, 2013; Pato et al., 2015; Lin and Li, 2019)).
@@
None of the particles of the standard model meet the requirements for explaining the numerous astrophysical observations and cosmological data that require the existence of dark matter. Thus candidate particles to constitute dark matter most often emerge from extensions of the standard model.
A much-studied class of such particles are WIMPs (weakly interacting massive particles), and many active experiments are attempting to reveal their existence. No such experiments, however, have so far obtained convincing results on the existence of DM particles even though very stringent constraints on the interaction cross section of dark matter with ordinary matter have been produced (see e.g. (Aprile et al., 2017)).
@@
Due to the very high densities ($10^{14}$$10^{15}\ \mathrm{g/cm^{3}}$) in their cores, the extreme surface gravity ($g\sim 10^{12}\,\mathrm{m/s^{2}}$) and their ages (up to $10^{10}\,\mathrm{yr}$), neutron stars are very good candidates where dark matter can be captured and can accumulate in sufficient amounts to influence their structural and evolutionary properties. Neutron stars can thus be used as natural detectors to identify DM and possibly unveil its nature (Bertone and Fairbairn, 2008; de Lavallaz and Fairbairn, 2010; Kouvaris and Tinyakov, 2010, 2011; Leung et al., 2011; Ciarcelluti and Sandin, 2011; Kouvaris et al., 2013; Kouvaris and Nielsen, 2015; Tolos and Schaffner-Bielich, 2015; Ellis et al., 2018; Deliyergiyev et al., 2019; Ivanytskyi et al., 2020; Del Popolo et al., 2020; Kain, 2020, 2021; Dengler et al., 2022; Shakeri and Karkevandi, 2024; Barbat et al., 2024; Mariani et al., 2024; Thakur et al., 2024).
@@
This work is organized as follows. In Section 2, we introduce the two-fluids Tolman-Oppenheimer-Volkoff (TOV) stellar structure equations. In Section 3, we discuss the equation of state (EOS) for OM and for DM used in the present work. In Section 4 we present results obtained by solving the two-fluids TOV equation and discuss the properties of DANS. In Section 5 we use observational data for properties of neutron stars to determine an upper limit to the dark matter particle mass. In Section 6, we study the stability of DANS equilibrium configurations with respect to small radial perturbations. Finally, in Section 7 we give a brief summary of our work.
2 Structure equations for DANS
Since the non-gravitational interaction between dark matter and ordinary matter is extremely small (e.g. (Bertone et al., 2005; Aprile et al., 2017)), it is possible to split the total energy-momentum tensor as the sum of the energy-momentum tensor of each of the two fluids (OM and DM) and to have covariant conservation for both of them. Accordingly, the equation of state of OM is independent on the state variables of DM and vice versa. In addition, it is assumed that each of the two fluids be a perfect fluid. Based on these assumptions, and further assuming a spherically symmetric and stationary distribution of OM and DM, the stellar structure equations in general relativity for DANS take the following form (see e.g. (Kain, 2020)), which generalizes the TOV equations to the case of two fluids interacting exclusively through gravity:
$\frac{dP_{j}}{dr}$ $=-G\,\frac{m_{tot}(r)\,\varepsilon_{j}(r)}{c^{2}r^{2}}\,\Bigg{(}1+\frac{P_{j}(r)}{\varepsilon_{j}(r)}\Bigg{)}$
×$\times\Bigg{(}1+\frac{4\pi r^{3}P_{tot}(r)}{c^{2}m_{tot}(r)}\Bigg{)}\,\Bigg{(}1-\frac{2\,G\,m_{tot}(r)}{c^{2}r}\Bigg{)}^{-1}$ (1)
@@ 2 Structure equations for DANS
2.1 Numerical integration
To solve the stellar structure equations (2) and (2) we need to specify the equation of state for the two fluids (see next section) and the appropriate boundary conditions at the center ($r=0$) and at the surface ($r=R_{j}$, $j=OM,DM$) of the matter distribution for each fluid (Kain, 2020):
@@ 2.1 Numerical integration
3 Equation of state
3.1 The equation of state of ordinary matter
Considering the existing uncertainties about the equation of state (EOS) of ordinary matter at high densities and considering that these uncertainties could masquerade the effects of dark matter on the observable properties of neutron stars, it is of paramount importance to employ an EOS for ordinary matter obtained using ab initio quantum many-body calculations well-grounded on modern theories of nuclear interactions.
In this work we model the OM fluid of a DANS as a uniform electric-charge-neutral fluid of neutrons, protons, electrons, and muons in equilibrium with respect to the weak interaction ($\beta$-stable nuclear matter) at zero temperature ($T=0$).
@@ 3.1 The equation of state of ordinary matter
Recently a new microscopic EOS for this system has been obtained in Ref. (Bombaci and Logoteta, 2018) (hereafter the BL EOS) for the zero temperature case, using the Brueckner-Hartree-Fock (BHF) quantum many-body approach (see (Bombaci and Logoteta, 2018) and references therein) starting from modern two-body and three-body nuclear interactions derived within chiral effective field theory (ChEFT) (e.g. (Machleidt and Entem, 2011; Hammer et al., 2020)). These chiral nuclear interactions reproduce with high accuracy the nucleon-nucleon (NN) scattering data and the experimental binding energies of light (A = 3, 4) atomic nuclei (Logoteta et al., 2016)).
This is a different approach with respect to other well-known microscopic EOS models for $\beta$-stable nuclear matter, as for example, the APR EOS (Akmal et al., 1998) or with respect to phenomenological EOS models derived, for example, from Skyrme nuclear interactions (see (Dutra et al., 2012) and references therein).
@@
The BL EOS reproduces the empirical saturation point (i.e. saturation density $n_{0}=0.16\pm 0.01~{}{\rm fm}^{-3}$, and energy per nucleon $E/A|_{n_{0}}=-16.0\pm 1.0~{}{\rm MeV}$) of symmetric nuclear matter, and other empirical properties (symmetry energy $E_{sym}$ and its slope parameter $L$, incompressibility) of nuclear matter at the saturation density $n_{0}$ (see Tab. 2 in (Bombaci and Logoteta, 2018)).
In particular the calculated $E_{sym}(n_{0})$ and $L$ for this EOS fulfill the bounds imposed by the unitary Fermi gas limit (Tews et al., 2017; Kievsky et al., 2018). As has been shown in ref. (Tews et al., 2017) several EOS models currently used in neutron star structure calculations, in supernova explosions and binary neutron star mergers simulations violate the unitary Fermi gas bounds. The calculated $E_{sym}(n_{0})$ and $L$ are also in agreement with the values $E_{sym}(n_{0})=(38.1\pm 4.7)\,\mathrm{MeV}$, $L=(106\pm 37)\,\mathrm{MeV}$ extracted in Ref. (Reed et al., 2021) from the measured neutron skin thickness of the 208Pb nucleus in the PREX-2 experiment (Adhikari et al., 2021).
@@
At supranuclear densities ($n>n_{0}$) the BL EOS model predicts a symmetry energy $E_{sym}(n)$ which is in very good agreement (see Fig. 3 in (Logoteta et al., 2016)) with the empirical constraints derived in (Danielewicz and Lee, 2014) using the excitation energies of isobaric analog states in nuclei. As is well known, the symmetry energy $E_{sym}(n)$, and particularly its density dependence is a crucial ingredient to determine the proton fraction in $\beta$-stable nuclear matter (Bombaci and Lombardo, 1991; Li et al., 2021; Zhang and Li, 2021; Bombaci, 2022) and ultimately it has an impact on the value of stellar radius and on the thickness of the crust (Lattimer and Prakash, 2001; Steiner, 2006) and on the possibility of having direct URCA processes (Lattimer et al., 1991) (see Section 4.2) and thus rapid cooling of neutron stars (Page et al., 2004). Further the BL EOS is consistent (see figure 2 in (Bombaci and Logoteta, 2018)) with the measured elliptic flow of matter in heavy-ion collisions experiments (Danielewicz et al., 2002).
@@
When computing static ordinary neutron star configurations the BL EOS (for the $\beta$-stable case) gives (Bombaci and Logoteta, 2018) a maximum mass $\rm{M}_{\rm max}=2.08\,\mathrm{M_{\odot}}$, with a corresponding central density $\rho_{c}=2.74\times 10^{15}\,\mathrm{g/cm^{3}}$ and radius $\rm{R}(M_{\rm max})=10.26\,\mathrm{km}$ and a quadrupolar tidal polarizability coefficient $\Lambda_{1.4}=385$ (for the $1.4\,\mathrm{M_{\odot}}$ neutron star (Logoteta and Bombaci, 2019)) compatible with the constraints derived from GW170817 (Abbott et al., 2017). Recently, the BL EOS has been extended (Logoteta et al., 2021) to finite temperature and to arbitrary proton fractions. This finite-temperature EOS model has been applied to numerical simulations of binary neutron star mergers (Bernuzzi et al., 2020; Endrizzi et al., 2018; Prakash et al., 2021). Finally, to model the (ordinary matter) neutron star crust (i.e. for nucleonic density $\leq 0.08$ fm-3) we have used the Baym–Pethick–Sutherland (Baym et al., 1971) and the Negele–Vautherin (Negele and Vautherin, 1973) EOS.
3.2 The equation of state of dark matter
@@
In the present work we consider non-self-annihilating fermionic DM. This so called fermionic asymmetric dark matter (ADM) (Kaplan et al., 2009; Zurek, 2014) carries a conserved charge which is analogous to the baryon number in the case of ordinary matter. We describe the ADM fluid as a non-interacting (ideal) gas of fermions with mass $m_{\chi}$ and spin $1/2$ at $T=0$. The corresponding EOS is well known (see e.g. (Shapiro and Teukolsky, 1983)) and is given by the following expressions
@@
4 DANS properties
For a fixed values of the dark matter particle mass $m_{\chi}$, a solution of the two-fluid TOV equations (2) and (2) is uniquely identified once we fix the central mass densities (central energy densities) $\rho_{cOM}$ and $\rho_{cDM}$ for the two fluids.
Figure 1: Mass-density profiles of ordinary matter (red) and dark matter (black) to the two-fluids TOV equation for $m_{\chi}=1\ \mathrm{GeV}$, $\rho_{cOM}=1.5\times 10^{15}\ \mathrm{g/cm^{3}}$ and different values of the dark matter central densities $\rho_{cDM}$. Each line-style correspond to a different choice of $\rho_{cDM}$: $(\rho_{cOM},0)$ (green continuous line), $(\rho_{cOM},10^{15}\ \mathrm{g/cm^{3}})$ (dotted lines), $(\rho_{cOM},4\times 10^{15}\ \mathrm{g/cm^{3}})$ (dash-dot lines), $(\rho_{cOM},2\times 10^{16}\ \mathrm{g/cm^{3}})$ (dashed lines).